Write the equation in standard form of the hyperbola with vertices (-3, 2) and (1, 2) and slope of asymptotes +-3/2.
Answer by lwsshak3(11628)

Write the equation in standard form of the hyperbola with vertices (-3, 2) and (1, 2) and slope of asymptotes +-3/2.
**
This is a hyperbola with horizontal transverse axis.
Its standard form of equation: (x-h)^2/a^2-(y-k)^2/b^2=1, (h,k)=(x,y) coordinates of center.
center:(-1,2)
length of horizontal transverse axis=4 (-3 to 1)=2a
a=2
a^2=4
..
given slope of asymptote=b/a=3/2
b=3a/2=3*2/2=3
b^2=9
..
Equation:
(x+1)^2/4-(y-2)^2/9=1